What is the largest possible value for the sum of five consecutive even numbers, if 10 and 12 are included amongst the five numbers?
(A) 90 (B) 50 (C) 40 (D) 60 (E) 70

Answer by stanbon(75887)

What is the largest possible value for the sum of five consecutive even numbers, if 10 and 12 are included amongst the five numbers?
--------------
10+12+14+16+18 = 2(28)+14 = 56+14 = 70
------------------------------
Cheers,
Stan H.
---------------------------

(A) 90 (B) 50 (C) 40 (D) 60 (E) 70